# Dynamics of ergotropy and environment-induced work

### v1

@@ I Introduction
In this work, we aim at exploring the environment-induced work, analytically relating it with ergotropy. We begin with the investigation of the dynamics of ergotropy in open systems under Markovian and non-Markovian evolutions. By expressing ergotropy for qubit states in terms of energy and coherence, we will split ergotropy in both coherent and incoherent contributions, obtaining the conditions for ergotropy sudden death and ergotropy freezing as a consequence of the system-bath interaction. These phenomena can be seen as analogous to the previously established sudden changes and freezing for classical and quantum correlations Maziero:09 ; Paula:13 , including entanglement Yu:09 . Then, we will establish a relationship between ergotropy variation and environment-induced work. More specifically, we will show that, for a constant Hamiltonian, ergotropy can be effectively changed by the environment-induced work, providing an interpretation of environment-induced work in terms of variation of ergotropy in the quantum system. The remaining contribution for the ergotropy variation is then given by the energy cost to transit between the initial and final passive states of the dynamics, with the passive state denoting the state with no energy available to be extracted. This result closely connects the environment-induced work with the ergotropy dynamics. We illustrate our relationship by looking at qubit states evolving under an amplitude damping process. For this example, we explicitly show that the ergotropy variation is an upper limit for the environment-induced work throughout the dynamics.
@@ II Ergotropy and its quantum dynamics
respectively. Note that $\mathcal{E}(C,0)=\mathcal{E}_{C}(C,0)=C$. According to Eq. (6), the dynamics of ergotropy will depend on the dynamics of both coherence $C$ and internal energy $U$. We can then determine explicit conditions for $U$ and $C$ that ensure peculiar behaviors for the ergotropy as a function of time. In particular, it is evident that
$\mathcal{E}(t)=\mathcal{E}_{0}\,\,$ iff $\,\,C(t)=C_{0}\,\,$ and $\,\,U(t)=U_{0}\,\,$  $(\forall\,t)$;
$\mathcal{E}(t)=0\,\,$ iff $\,\,C(t)=0\,\,$ and $\,\,U(t)\leq 0\,\,$ $(\forall\,t)$,
where $\mathcal{E}_{0}=\mathcal{E}(0)$, $C_{0}=C(0)$, and $U_{0}=U(0)$. The properties (a) and (b) reveal that the dynamics of ergotropy can exhibit freezing and sudden death effects, respectively. In order to explore these two phenomena, we start by defining the dynamics of a quantum system interacting with an environment in terms of Kraus representation as $\rho(t)=\sum_{i}K_{i}(t)\rho(0)K_{i}(t)^{\dagger}$, where the Kraus operators $\{K_{i}(t)\}$ satisfy the completeness relation $\sum_{i}K_{i}(t)^{\dagger}K_{i}(t)=I$ and allows the environment characterization in both Markovian and non-Markovian regimes.
@@ II.1 Freezing
To illustrate this phenomenon, we show in Fig. 1 the ergotropy $\mathcal{E}$ as a function of $\gamma t$ for $C_{0}=0.5$ and $U_{0}=0.5$. For the sake of comparison, its coherent [Eq. (7)] and incoherent [Eq. (8)] contributions are also included on the figure. In both non-Markovian and Markovian regimes, notice that the incoherent component freezes for all $\gamma t$, while the coherent part exhibits a monotonic decay. In terms of the total ergotropy, the contribution of its coherent component allows the ergotropy to decay until it achieves a steady state. Additionally, it is clear to see in Fig. 1 a delayed ergotropy decay in the non-Markovian regime in comparison with the Markovian regime.
@@ II.2 Sudden death
In the Markovian regime, $\Gamma\rightarrow\infty$, the emergence of eternal death is determined by $t_{sd}=t_{1}=\gamma^{-1}\ln(1+U_{0})$. In this limit, there are no temporary collapses and revivals.
Figure 2 shows the ergotropy, as well as its incoherent and coherent parts, as functions of $\gamma t$ for the initial conditions $C_{0}=0$ and $U_{0}=0.5$. Remarkably, the incoherent component exhibits a non-monotonic (monotonic) decay until the eternal death time $t_{sd}\approx 297/\gamma$ ($t_{sd}\approx 0.405/\gamma$) for the non-Markovian (Markovian) regime, with $\Gamma=0.001\gamma$ ($\Gamma\rightarrow\infty$). Besides, notice that the coherent part contributes to the inhibition of this phenomenon.
@@ III Environment-induced work
$dU=\delta\mathcal{Q}+\delta\mathcal{W},$ (16)
where $dU$ is the infinitesimal internal energy change, $\delta\mathcal{Q}$ is the infinitesimal heat exchanged in the process, and $\delta\mathcal{W}$ is the infinitesimal work performed by (or on) the system, with $U=\text{Tr}(H\rho)$, $\delta\mathcal{Q}=\text{Tr}(H\,d\rho)$, and $\delta\mathcal{W}=\text{Tr}(dH\,\rho)$. As an alternative formulation, we can modify the definition of $\delta\mathcal{Q}$ so that heat is directly linked with the entropy variation. In this entropy-based framework for quantum thermodynamics adolfo , heat and work are defined through $\delta Q=\delta\mathcal{Q}-\delta W^{*}$ and $\delta W=\delta\mathcal{W}+\delta W^{*}$, so that the first law of thermodynamics is
@@
$\delta W^{*}=\sum_{n}r_{n}\left(\left\langle r_{n}\right|H\,d\left|r_{n}\right\rangle+h.c.\right).$ (18)
Notice that the first law of thermodynamics is preserved, with the internal energy infinitesimal $dU$ kept unchanged due to the new definitions of heat and work.
A utility for the definition of heat $Q$ as a witness of non-Markovianity for unital quantum maps has been recently provided john . Now, we will provide an operational meaning for $W^{*}$ in terms of ergotropy variation. We consider a quantum system described by an initial density operator $\rho(t=0)\to\rho_{0}$ and governed by a constant Hamiltonian $H$. In this scenario, the conventional work $\mathcal{W}$ is null. The system interacts with an external environment and is taken to a final density operator $\rho(t_{c})\to\rho_{c}$ at a specific characteristic time $t_{c}$ such that the total heat exchanged with the environment is vanishing. For this effective adiabatic process, the environment-induced work is the only contribution to the energy balance in the first law of thermodynamics, i.e.,
@@
The contribution $\Delta U_{\pi}(\rho_{c})$ for the ergotropy can be interpreted as the energy cost to transit between the initial and final passive states $\rho_{0\pi}$ and $\rho_{c\pi}$, respectively. We observe that Eq. (24) agrees with the discussion about the energetics of the ergotropy in Ref. binder , with $\Delta U_{\pi}$ defined there as an operational heat. Here, we can then directly connect the environment-induced work $W^{*}$ with the variation of ergotropy for constant Hamiltonians, reinforcing the interpretation of $W^{*}$ as an effective work extracted due to the system-environment interaction. For 2-level systems, the Hamiltonian is given by $H=-\sigma_{z}$ and the density matrix, in terms of the Bloch sphere, can be written as $\rho(t)=\left(I+\vec{r}(t)\cdot\vec{\sigma}\right)/2$, where $\vec{r}(t)=[x(t),y(t),z(t)]$ is the Bloch vector and $I,\vec{\sigma}$ are Pauli operators. Since $\delta Q=(U/r)dr$ with $U(t)=-z(t)$  john , the characteristic thermal adiabatic time $t_{c}$ can be obtained through the condition
@@
Thus, the PD process discussed in Sec. II (a) is unable to extract the available quantum resource $\mathcal{E}_{C}$ through environment-induced work. In other words, there is no effective adiabatic process with $W^{*}(t_{c})\neq 0$ for PD channels. On the other hand, the extraction is possible for dissipative quantum processes such as the AD channel described in Sec.II (b).
Here, we will illustrate the connection between $W^{*}$ and $\Delta\mathcal{E}$ by considering the paradigmatic model of the decay of an excited state of a 2-level atom interacting with an environment by spontaneous emission Nielsen-Chuang (a Markovian AD process). The spontaneous emission process is governed by the Markovian master equation
$\dot{\rho}(t)=-i\left[\sigma_{z},\rho(t)\right]+\gamma\left[\sigma^{-}\rho(t)\sigma^{+}-\frac{1}{2}\left\{\sigma^{+}\sigma^{-},\rho(t)\right\}\right],$ (29)
where $\gamma$ is the dissipation rate of spontaneous emission and $\sigma^{+},\sigma^{-}$ are the raising and lowering operators for a 2-level atom. The solution of Eq. (29) with an arbitrary initial state $\vec{r}_{0}=[x_{0},y_{0},z_{0}]$ is given by $\vec{r}(t)=[x(t),y(t),z(t)]$ with
@@
According to this solution and Eq. (25), the dimensionless characteristic adiabatic time $\tau_{c}=\gamma t_{c}$ is a function only of the initial parameters $r_{0}=(x_{0}^{2}+y_{0}^{2}+z_{0}^{2})^{1/2}$ and $\theta_{0}=\text{arccos}[z_{0}/r_{0}]$, where $0\leq r_{0}\leq 1$ and $0\leq\theta_{0}\leq\pi$. We numerically investigate the adiabatic time for arbitrary initial states $\vec{r}_{0}=(r_{0},\theta_{0})$ in Fig. 3. Notice that $\tau_{c}$ is not negligible for initial states close to $\theta_{0}=\pi/2$ or $r_{0}=1$, mainly for $\theta_{0}\geq\pi/2$ (north hemisphere), where $\Delta\mathcal{E}=\Delta\mathcal{E}_{C}$. We also study the energy cost $\Delta U_{\pi}$ as a function of $r_{0}$ and $\theta_{0}$. This is exhibited in Fig. 4. Notice that $\Delta U_{\pi}\geq 0$ for all initial states. Therefore, from Eq. (24), the work $W^{*}(t_{c})$ performed on the system by the environment is not greater (in absolute value) than the ergotropy variation $\Delta\mathcal{E}(t_{c})$. This result implies that the environment cannot provide more energy to the system than it can be extracted via the definition of extractable work through a variation of ergotropy. Finally, we plot $\Delta\mathcal{E}(t_{c})$ and $W^{*}(t_{c})$ for two particular types of families of initial states: a family of mixed states located on the equatorial plane of the Bloch sphere (see Fig. 5) and a family of pure states located on the upper surface of the Bloch sphere (see Fig. 6). As a by-product, we can also determine the energy cost associated with the remaining contribution $\Delta U_{\pi}(t_{c})$ for the ergotropy variation, as shown in the inset of Fig. 5 and Fig. 6. In theses cases, it is evident that $W^{*}(t_{c})/\Delta\mathcal{E}(t_{c})\leq 1$, where $\Delta\mathcal{E}(t_{c})=\Delta\mathcal{E}_{C}(t_{c})$.
We can also consider a more general dynamics such as the non-Markovian case. This can be analyzed using the same physical process as in Eq. (29), but with a time-dependent dissipation rate $\gamma(t)$ just like discussed in the sudden death section. For each initial state, instead of only a single $t_{c}$, a set of characteristic times emerges, denoted by $\{t_{nc}\}$. By looking at these characteristic times, we can show that the largest $t_{nc}$ in the set, for each state, turns out to present results very close to the Markovian case, as already shown in Fig. 5 and Fig. 6.
@@
Figure 5: (Color online) Environment-induced work $W^{*}$, ergotropy variation $\Delta\mathcal{E}$, and passive energy cost $\Delta U_{\pi}$ (inset) as functions of $r_{0}$ ($0\leq r_{0}\leq 1$) for $\theta_{0}=\pi/2$.
@@
Figure 6: (Color online) Environment-induced work $W^{*}$, ergotropy variation $\Delta\mathcal{E}$, and passive energy cost $\Delta U_{\pi}$ (inset) as functions of $\theta_{0}$ ($20\leq\theta_{0}\leq\pi/$ 2) for $r_{0}=1$.

### v2

# Qubit dynamics of ergotropy and environment-induced work

@@ I Introduction
In this work, we aim at exploring the environment-induced work, analytically relating it with ergotropy. We begin with the investigation of the dynamics of ergotropy in open systems under Markovian and non-Markovian evolutions. By expressing ergotropy for qubit states in terms of energy and coherence, we will split ergotropy in both coherent and incoherent contributions, obtaining the conditions for ergotropy sudden death and ergotropy freezing as a consequence of the system-bath interaction. These phenomena can be seen as analogous to the previously established sudden changes and freezing for classical and quantum correlations Maziero:09 ; Paula:13 , including entanglement Yu:09 . Then, we will establish a relationship between ergotropy variation and environment-induced work. More specifically, we will show that, for a constant Hamiltonian, ergotropy can be effectively changed by the environment-induced work, providing an interpretation of environment-induced work in terms of variation of ergotropy in the quantum system. The remaining contribution for the ergotropy variation is then given by the energy cost to transit between the initial and final passive states of the dynamics, with the passive state denoting the state with no energy available to be extracted. This result closely connects the environment-induced work with the ergotropy dynamics.
We illustrate our relationship by looking at qubit states evolving under an amplitude damping process. For this example, we explicitly show that the ergotropy variation is an upper limit for the environment-induced work throughout the dynamics.
@@ II Ergotropy and its quantum dynamics
respectively. Note that $\mathcal{E}(C,0)=\mathcal{E}_{C}(C,0)=C$. According to Eq. (6), the dynamics of ergotropy will depend on the dynamics of both coherence $C$ and internal energy $U$. We can then determine explicit conditions for $U$ and $C$ that ensure peculiar behaviors for the ergotropy as a function of time. In particular, it is evident that (a) $\mathcal{E}(t)=\mathcal{E}_{0}\,\,$ iff $\,\,C(t)=C_{0}\,\,$ and $\,\,U(t)=U_{0}\,\,$  $(\forall\,t)$; (b) $\mathcal{E}(t)=0\,\,$ iff $\,\,C(t)=0\,\,$ and $\,\,U(t)\leq 0\,\,$ $(\forall\,t)$, where $\mathcal{E}_{0}=\mathcal{E}(0)$, $C_{0}=C(0)$, and $U_{0}=U(0)$. The properties (a) and (b) reveal that the dynamics of ergotropy can exhibit freezing and sudden death effects, respectively. In order to explore these two phenomena, we start by defining the dynamics of a quantum system interacting with an environment in terms of Kraus representation as $\rho(t)=\sum_{i}K_{i}(t)\rho(0)K_{i}(t)^{\dagger}$, where the Kraus operators $\{K_{i}(t)\}$ satisfy the completeness relation $\sum_{i}K_{i}(t)^{\dagger}K_{i}(t)=I$ and allows the environment characterization in both Markovian and non-Markovian regimes.
@@ II.1 Freezing
As a consequence of Eqs. (7) and (11), $\mathcal{E}_{I}(t)=2U_{0}$ $\forall\,t$ iff $U_{0}>0$ (for all $C_{0}$). To illustrate this phenomenon, we show $\mathcal{E}$, $\mathcal{E}_{I}$, and $\mathcal{E}_{C}$ in Fig. 1 as functions of $\gamma t$ for $C_{0}=0.5$ and $U_{0}=0.5$, where the initial state is taken as $\rho_{00}=0.25$, $\rho_{11}=1-\rho_{00}$, and $|\rho_{01}|=0.25$. In both non-Markovian and Markovian regimes, notice that the incoherent component freezes for all $\gamma t$, while the coherent part exhibits a monotonic decay. In terms of the total ergotropy, the contribution of its coherent component allows the ergotropy to decay until it achieves a steady state. Additionally, it is clear to see in Fig. 1 a delayed ergotropy decay in the non-Markovian regime in comparison with the Markovian regime.
@@ II.2 Sudden death
Consequently, $\mathcal{E}_{I}(t)=0$ $\forall\,t\geq t_{sd}$ iff $U_{0}>0$ (for all $C_{0}$). In the Markovian regime, $\Gamma\rightarrow\infty$, the emergence of eternal death is
@@
determined by $t_{sd}=t_{1}=\gamma^{-1}\ln(1+U_{0})$. In this limit, there are no temporary collapses and revivals.
Figure 2 shows the ergotropy, as well as its incoherent and coherent parts, as functions of $\gamma t$ for the initial conditions $C_{0}=0.5$, $U_{0}=0.5$, and initial state given by $\rho_{00}=0.25$, $\rho_{11}=1-\rho_{00}$, and $|\rho_{01}|=0.25$. Remarkably, the incoherent component exhibits a non-monotonic (monotonic) decay until the eternal death time $t_{sd}\approx 297/\gamma$ ($t_{sd}\approx 0.405/\gamma$) for the non-Markovian (Markovian) regime, with $\Gamma=0.001\gamma$ ($\Gamma\rightarrow\infty$). Besides, notice that the coherent part contributes to the inhibition of this phenomenon.
@@ III Environment-induced work
$dU=\delta\tilde{Q}+\delta\tilde{W},$ (16)
where $dU$ is the infinitesimal internal energy change, $\delta\tilde{Q}$ is the infinitesimal heat exchanged in the process, and $\delta\tilde{W}$ is the infinitesimal work performed by (or on) the system, with $U=\text{Tr}(H\rho)$, $\delta\tilde{Q}=\text{Tr}(H\,d\rho)$, and $\delta\tilde{W}=\text{Tr}(dH\,\rho)$. As an alternative formulation, we can modify the definition of $\delta\tilde{Q}$ so that heat is directly linked with the entropy variation. In this entropy-based framework for quantum thermodynamics adolfo , heat and work are defined through $\delta Q=\delta\tilde{Q}-\delta W^{*}$ and $\delta W=\delta\tilde{W}+\delta W^{*}$, so that the first law of thermodynamics is
@@
$\delta W^{*}=\sum_{n}r_{n}\left(\left\langle r_{n}\right|H\,d\left|r_{n}\right\rangle+h.c.\right),$ (18)
where |$\ket{r_{n}}$ represents an eigenvector of the density operator $\rho$ and $r_{n}$ the corresponding eigenvalue. Notice that the first law of thermodynamics is preserved, with the internal energy infinitesimal $dU$ kept unchanged due to the new definitions of heat and work. From this point on, unless stated otherwise, heat and work will refer to the entropy-based formulation of quantum thermodynamics.
A utility for the definition of heat $Q$ as a witness of non-Markovianity for unital quantum maps has been recently provided john . Now, we will provide an operational meaning for $W^{*}$ in terms of ergotropy variation. We consider a quantum system described by an initial density operator $\rho(t=0)\to\rho_{0}$ and governed by a constant Hamiltonian $H$. In this scenario, the conventional work $\tilde{W}$ is null. The system interacts with an external environment and is taken to a final density operator $\rho(t_{c})\to\rho_{c}$ at a specific characteristic time $t_{c}$ such that the total heat $Q$ exchanged with the environment is vanishing. For this effective adiabatic process, the environment-induced work is the only contribution to the energy balance in the first law of thermodynamics, i.e.,
@@
The contribution $\Delta U_{\pi}(\rho_{c})$ for the ergotropy can be interpreted as the energy cost to transit between the initial and final passive states $\rho_{0\pi}$ and $\rho_{c\pi}$, respectively. We observe that Eq. (24) agrees with the discussion about the energetics of the ergotropy in Ref. binder , with $\Delta U_{\pi}$ defined there as an operational heat. Here, we can then directly connect the environment-induced work $W^{*}$ with the variation of ergotropy for constant Hamiltonians, reinforcing the interpretation of $W^{*}$ as an effective work extracted due to the system-environment interaction. For 2-level systems, the Hamiltonian is given by $H=-\sigma_{z}$ and the density matrix, in terms of the Bloch sphere, can be written as $\rho(t)=\left(I+\vec{r}(t)\cdot\vec{\sigma}\right)/2$, where $\vec{r}(t)=[x(t),y(t),z(t)]$ is the Bloch vector and $I,\vec{\sigma}$ are Pauli operators. In this case, coherence and internal energy are given by $C(t)=\sqrt{x(t)^{2}+y(t)^{2}}$ and $U(t)=-z(t)$, respectively. Since $\delta Q=(U/r)dr$  john , the
characteristic adiabatic time $t_{c}$ can be obtained through the condition
@@
Thus, the PD process discussed in Sec. II.1 is unable to extract the available quantum resource $\mathcal{E}_{C}$ through environment-induced work. In other words, there is no effective adiabatic process with $W^{*}(t_{c})\neq 0$ for PD channels. On the other hand, the extraction is possible for dissipative quantum processes such as the AD channel described in Sec. II.2.
Here, we will illustrate the connection between $W^{*}$ and $\Delta\mathcal{E}$ by considering the paradigmatic model of the decay of an excited state of a 2-level atom interacting with an environment by spontaneous emission Nielsen-Chuang ; Breuer:09 (a Markovian AD process). The spontaneous emission process is governed by the Markovian master equation
$\dot{\rho}(t)=i\left[\sigma_{z},\rho(t)\right]+\gamma\left[\sigma^{-}\rho(t)\sigma^{+}-\frac{1}{2}\left\{\sigma^{+}\sigma^{-},\rho(t)\right\}\right],$ (29)
where $\gamma$ is the dissipation rate of spontaneous emission and $\sigma^{+}=(\sigma_{x}-i\sigma_{y})/2$ and $\sigma^{-}=(\sigma_{x}+i\sigma_{y})/2$ are the raising and lowering operators for a 2-level atom. Notice that the ground state of $H$ is the computational state $|0\rangle$ in the north pole of the Bloch sphere, which is the expected long time limit after energy loss in the spontaneous emission dynamics. The solution of Eq. (29) with an arbitrary initial state $\vec{r}_{0}=[x_{0},y_{0},z_{0}]$ is given by $\vec{r}(t)=[x(t),y(t),z(t)]$ with
@@
According to this solution and Eq. (25), the dimensionless characteristic adiabatic time $\tau_{c}=\gamma t_{c}$ is a function only of the initial parameters $r_{0}=(x_{0}^{2}+y_{0}^{2}+z_{0}^{2})^{1/2}$ and $\theta_{0}=\text{arccos}[z_{0}/r_{0}]$, where $0\leq r_{0}\leq 1$ and $0\leq\theta_{0}\leq\pi$. We numerically investigate
@@
Figure 5: (Color online) Environment-induced work in the Markovian $W^{*}$ and non-Markovian $W^{*}_{n}$ regimes, as well as ergotropy variation in the Markovian $\Delta\mathcal{E}$ and non-Markovian $\Delta\mathcal{E}_{n}$ regimes, as functions of $r_{0}$ ($0\leq r_{0}\leq 1$) for $\theta_{0}=\pi/2$. For the non-Markovian dynamics, we adopted $\Gamma=0.01\gamma$. Inset: Passive energy cost for both Markovian $\Delta U_{\pi}$ and non-Markovian $\Delta U_{\pi n}$ regimes.
@@
Figure 6: (Color online) Environment-induced work in the Markovian $W^{*}$ and non-Markovian $W^{*}_{n}$ regimes, as well as ergotropy variation in the Markovian $\Delta\mathcal{E}$ and non-Markovian $\Delta\mathcal{E}_{n}$ regimes, as functions of $\theta_{0}$ ($20\leq\theta_{0}\leq\pi/$ 2) for $r_{0}=1$. For the non-Markovian dynamics, we adopted $\Gamma=0.01\gamma$. Inset: Passive energy cost for both Markovian $\Delta U_{\pi}$ and non-Markovian $\Delta U_{\pi n}$ regimes.
the characteristic adiabatic time for arbitrary initial states $\vec{r}_{0}=(r_{0},\theta_{0})$ in Fig. 3. Notice that $\tau_{c}$ is not negligible for initial states close to $\theta_{0}=\pi/2$ or $r_{0}=1$, mainly for $\theta_{0}\geq\pi/2$ (north hemisphere), where $\Delta\mathcal{E}=\Delta\mathcal{E}_{C}$. We also study the energy cost $\Delta U_{\pi}$ as a function of $r_{0}$ and $\theta_{0}$. This is exhibited in Fig. 4. Notice that $\Delta U_{\pi}\geq 0$ for all initial states. Therefore, from Eq. (24), the work $W^{*}(t_{c})$ performed on the system by the environment is not greater (in absolute value) than the ergotropy variation $\Delta\mathcal{E}(t_{c})$. This result implies that the environment cannot provide more energy to the system than it can be extracted via the definition of extractable work through a variation of ergotropy. Finally, we plot $\Delta\mathcal{E}(t_{c})$ and $W^{*}(t_{c})$ for two particular types of families of initial states: a family of mixed states located on the equatorial plane of the Bloch sphere (see Fig. 5) and a family of pure states located on the upper surface of the Bloch sphere (see Fig. 6), where $t_{c}\neq 0$. As a by-product, we can also determine the energy cost associated with the remaining contribution $\Delta U_{\pi}(t_{c})$ for the ergotropy variation, as shown in the inset of Fig. 5 and Fig. 6. In theses cases, it is evident that $W^{*}(t_{c})/\Delta\mathcal{E}(t_{c})\leq 1$, where $\Delta\mathcal{E}(t_{c})=\Delta\mathcal{E}_{C}(t_{c})$. Notice also that the singular behavior of the dimensionless characteristic time shown in Fig. 3 is not manifested in Figs. 5 and 6. This occurs because we have kept the dynamics in the north hemisphere of the Bloch sphere, with $\theta_{0}\leq\pi/2$.
We can also consider a more general dynamics, such as the non-Markovian case. This can be analyzed using the physical process as in Eq. (12), from which Eq. (29) follows as a Markovian limit. For each initial state, instead of only a single $t_{c}$, a set of characteristic times emerges, which is denoted by $\{t_{nc}\}$. By looking at these characteristic times, we show that, by taking the largest $t_{nc}$ for each state, we obtain results very close to the Markovian case, as shown in Fig. 5 and Fig. 6. This behavior holds independently of the ratio $\Gamma/\gamma$.